# Fast Frank–Wolfe Algorithms with Adaptive Bregman Step-Size for Weakly Convex Functions

### v3

@@ C.1 Sublinear Convergence
where the first inequality holds because of Lemma 2.4 and the last inequality holds because of Lemma 3.2. Subtracting $f^{*}$ on both sides, using $D_{\phi}(v_{t},x_{t})\leq D^{2}$, and rearranging leads to
@@
If $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu=1$ in (C.1), we have $f(x_{t})-f^{*}\leq 4LD^{2}/(t+2)$, which is the same as a sublinear convergence rate of the classical FW algorithm.
C.2 Proof of Theorem 4.2
Because Algorithm 2 is well-defined (see Remark 3.5), the convergence result of the FW algorithm using the adaptive step-size strategy (Algorithm 2) is essentially the same as the one that uses Bregman short steps (Theorems 4.2 and 5.2); up to small errors arising from the approximation whose precise analysis we skip for the sake of brevity.
Theorem C.2 (Linear convergence FW algorithm with short step-size or adaptive step-size).
Suppose that Assumptions 3.1 and 4.1 hold. Let $\nu>0$ and let $f$ satisfy the HEB condition with $q\geq 1+\nu$ and $\mu>0$ and $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ be the diameter of $P$. Assume that there exists a minimizer $x^{*}\in\operatorname{int}P$, i.e., there exists an $r>0$ with $B(x^{*},r)\subset P$. Consider the iterates of Algorithm 1 with $\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{1/\nu},1\right\}$ or Algorithm 2. Then, it holds that, for all $t\geq 1$,
$f(x_{t})-f^{*}\leq\begin{cases}h_{t}\leq\max\left\{\frac{1}{1+\nu},1-\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{c^{1+1/\nu}D^{2/\nu}}\right\}^{t-1}LD^{2}&\text{if}\ q=1+\nu,\\ \frac{LD^{2}}{(1+\nu)^{t-1}}&\text{if}\ 1\leq t\leq t_{0},q>1+\nu,\\ \frac{(L(1+\nu)(c/r)^{1+\nu}D^{2})^{q/(q-1-\nu)}}{\left(1+\frac{1-\nu}{2(1+\nu)}(t-t_{0})\right)^{\nu q/(q-1-\nu)}}=\mathcal{O}(1/t^{\nu q/(q-1-\nu)})&\text{if}\ t\geq t_{0},q>1+\nu,\end{cases}$
where $c=(q/\mu)^{1/q}$ and
$t_{0}:=\max\left\{\left\lfloor\log_{\frac{1}{1+\nu}}\left(\frac{(L(1+\nu(c/r)^{1+\nu}D^{2})^{q/(q-1-\nu)}}{LD^{2}}\right)\right\rfloor+2,1\right\}.$
Proof.
Let $h_{t}:=f(x_{t})-f^{*}$. Using Lemma 3.4, we have
@@
where the second inequality holds from $\langle f(x_{t}),x_{t}-v_{t}\rangle\geq r\|\nabla f(x_{t})\|$ in Proposition B.7, and the last inequality holds because $f$ satisfies the $c$-Łojasiewicz inequality with $c=(q/\mu)^{1/q}$ (see Remark B.4), and $M:=(L(1+\nu))^{1/\nu}$.
(ii) $\gamma_{t}=1$: Using (3.1), we have
@@
This inequality and the initial bound $f(x_{1})-f^{*}\leq LD^{2}$ due to Lemma A.4 imply
@@
Using $f(x_{1})-f^{*}\leq LD^{2}$ and Lemma B.11 with $c_{0}=LD^{2}$, $c_{1}=\frac{\nu}{1+\nu}$, $c_{2}=\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}$, and $\theta_{0}=\frac{q-1-\nu}{\nu q}>0$ from $q>1+\nu$, we have the claim. ∎
C.3 Proof of Theorem 4.4
Theorem C.3 (Linear convergence of the away-step FW algorithm).
Suppose that Assumptions 3.14.1, and 4.3 hold. Let $\nu>0$ and let $P\subset\mathbb{R}^{n}$ be a polytope and $f$ satisfy the HEB condition with $q>1+\nu$ or $(\nu,q)=(1,2)$. The convergence rate of Algorithm 3 is linear: for all $t\geq 1$
where $c:=(q/\mu)^{1/q}$ $D$ and $\delta$ are the diameter and the pyramidal width of the polytope $P$, respectively, and
$t_{0}:=\max\left\{\left\lfloor\log_{\frac{1}{1+\nu}}\frac{(L(1+\nu)D^{2}/((\mu/q)^{1/q}\delta/2)^{1+\nu})^{q/(q-1-\nu)}}{LD^{2}}\right\rfloor+2,1\right\}.$
Proof.
By using the induced guarantee on the primal gap via Lemma B.10, we have
$h_{t}=f(x_{t})-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{\langle\nabla f(x_{t}),v_{t}^{\textup{A}}-v_{t}^{\textup{FW}}\rangle}{\delta}\right)^{\frac{q}{q-1}}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{2\langle\nabla f(x_{t}),d_{t}\rangle}{\delta}\right)^{\frac{q}{q-1}},$ (C.2)
@@
where the first inequality holds due to Lemma 3.4 with $\gamma_{t,\max}=1$ (Frank–Wolfe steps) and $\gamma_{t,\max}=\frac{\lambda_{v_{t}^{\textup{A}},t}}{1-\lambda_{v_{t}^{\textup{A}},t}}$ (away steps), and the second inequality holds because of (C.2) and $\langle\nabla f(x_{t}),d_{t}\rangle\geq\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\geq h_{t}$.
@@
In $q>1+\nu$, we have $h_{t}-h_{t+1}\geq\min\left\{\frac{\nu}{1+\nu},\frac{\nu}{1+\nu}\frac{((\mu/q)^{1/q}\delta/2)^{\frac{1+\nu}{\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}h_{t}^{\frac{q-1-\nu}{\nu q}}\right\}\cdot h_{t}$ for at least half of the iterations (Frank–Wolfe steps) and $h_{t+1}\leq h_{t}$ for the rest (away steps). The initial bound $f(x_{1})-f^{*}\leq LD^{2}$ holds generally for the Frank–Wolfe algorithm (see the proof of Theorem C.1 and [Braun2022-ew, Remark 2.4]). We use Lemma B.11 with $c_{0}=LD^{2}$, $c_{1}=\frac{\nu}{1+\nu}$, $c_{2}=c_{1}\frac{((\mu/q)^{1/q}\delta/2)^{\frac{1+\nu}{\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}$, and $\theta_{0}=\frac{q-1-\nu}{\nu q}>0$ from $q>1+\nu$ and obtain the claim. ∎
Remark C.4 (Compatibility of parameters).
@@
$\frac{((\mu/q)^{1/q}\delta/2)^{\frac{1+\nu}{\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}=\left(\frac{\mu}{L}\right)^{1/\nu}\left(\frac{\delta}{D}\right)^{\frac{1+\nu}{\nu}}\frac{1}{2^{\frac{1+\nu}{\nu}}(1+\nu)^{2/\nu}D^{(1-\nu)/\nu}},$ (C.3)
which might be greater than 1 because $\frac{1}{D^{(1-\nu)/\nu}}\geq 1$ if $D<1$ and $\nu$ are small enough. In order to make (C.3) smaller than 1, $\nu$ should be 1, i.e., $(\nu,q)=(1,2)$. When $\nu=1$ and $q\neq 2$, it reduces to the $q>1+\nu$ case.
Appendix D Proof of Convergence Analysis for Nonconvex Optimization
D.1 Global Convergence
We show that Algorithm 1 with $\gamma_{t}=\gamma:=1/(T+1)^{\frac{1}{1+\nu}}$ globally converges to a stationary point, where $T\in\mathbb{N}$ is the number of iterations. Its proof is inspired by [Lacoste-Julien2016-mj, Theorem 1] and [Pokutta2024-qu, Theorem 4.7] and identical to those for the case when $\phi=\frac{1}{2}\|\cdot\|^{2}$.
Theorem D.1 (Global sublinear convergence for nonconvex optimization).
@@ D.1 Global Convergence
Substituting $x_{t+1}$ for $x_{+}$ and $x_{t}$ for $x$ in Lemma 2.4, we have the following inequality:
@@
which is the same rate as [Lacoste-Julien2016-mj, Theorem 1].
D.2 Local Sublinear Convergence
@@
Now we show sublinear convergence with $\gamma_{t}=\frac{2}{2+t}$. The proof is a modified version of Theorem C.1.
Theorem D.3 (Local sublinear convergence).
@@
Using Lemma 2.4, we have
@@
where the last inequality holds because of (3.2) in Lemma 3.3. Subtracting $f^{*}$ and rearranging provides
$f(x_{t+1})-f^{*}$ $\leq(1-\gamma_{t})(f(x_{t})-f^{*})+\frac{\rho\gamma_{t}}{2}\|x_{t}-x^{*}\|^{2}+L\gamma_{t}^{1+\nu}D_{\phi}(v_{t},x_{t})$ (D.3)
$\leq\left(1-\left(1-\frac{\rho}{\mu}\right)\gamma_{t}\right)(f(x_{t})-f^{*})+L\gamma_{t}^{1+\nu}D^{2},$ (D.4)
where the last inequality holds due to the quadratic growth condition. For $t=0$, using (D.3), we have
$f(x_{1})-f^{*}\leq\frac{\rho}{2}\|x_{0}-x^{*}\|^{2}+LD_{\phi}(v_{0},x_{0})\leq\left(\frac{\rho}{\sigma}+L\right)D^{2}\leq 2LD^{2}\leq\frac{2\mu}{\rho}LD^{2},$ (D.5)
where the second inequality holds because $\phi$ is $\sigma$-strongly convex and the last inequality holds because of $3<6/(2-\nu)\leq 2\mu/\rho$. Now we consider $t\geq 1$. Using (D.4) and $\gamma_{t}=\frac{2}{2+t}$, we have
@@
where the last inequality holds because of $(t+3)^{\nu}(t+3\rho/\mu)\leq(t+3)^{\nu}(t+2-\nu)\leq(t+2)^{1+\nu}$ from (D.1). ∎
@@
Note that Theorem D.3 does not require knowledge of the number of iterations $T$ ahead of time compared to Theorem D.1. We stress, nonetheless, that the latter can also be adjusted using a different step-size strategy to obtain an any-time guarantee; see [Braun2022-ew] for details for the standard Euclidean case, which can be generalized to our setup. Moreover, we can apply Theorem D.3 to the Euclidean FW algorithm, i.e., $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu=1$. Using (D.2), $D_{\textup{Euc}}:=\sup_{x,y\in P}\|x-y\|^{2}$, and $D_{\textup{Euc}}=\sqrt{2}D$, we obtain
@@
Theorem D.3 requires $\rho/\sigma\leq L$ and $\frac{3\rho}{\mu}\leq 2-\nu$. These assumptions are easy to satisfy.
Example D.4 (Example 3.1 in the arXiv version of [Liao2024-vd]).
@@
The function $f$ is not convex but $\rho$-weakly convex with $\rho=2$. A global optimal solution of $f$ is $x^{*}=-1$ and its value is $f(x^{*})=0$. Moreover, $f$ satisfies the quadratic growth condition with $0<\mu\leq 6$. Because $f$ is a quadratic function, we have $L\geq 6$. It holds that $2=\rho\leq L$ and $1=\frac{3\rho}{\mu}\leq 1<2-\nu$ (set $\mu=6$). Therefore, the assumption of Theorem D.3 holds.
In order to verify $\frac{3\rho}{\mu}\leq 2-\nu$, it is easier to examine the sufficient condition $\frac{3\rho}{\mu}\leq 1$ instead because the exact value of $\nu$ is difficult to estimate. On the other hand, without loss of generality, we can assume $\sigma=1$. When $\sigma>1$, $\rho/\sigma<\rho\leq L$ holds. When $\sigma<1$, we can use $\phi_{1}=\phi+\frac{1-\sigma}{2}\|\cdot\|^{2}$, which is 1-strongly convex. When $\phi$ is convex but not strongly convex, we can use $\phi_{2}=\phi+\frac{1}{2}\|\cdot\|^{2}$. Therefore, it suffices to verify $\rho/\sigma\leq\rho\leq L$, which often holds (for example, see an example of phase retrieval in Section 6.2). Next we establish local linear convergence with the short step step-size, i.e., $\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}},1\right\}$.
D.3 Proof of Theorem 5.2
Theorem D.5 (Local linear convergence of FW algorithm with short step-size or adaptive step-size).
Suppose that Assumptions 3.1 and 5.1 hold. Let $\nu\in(0,1]$ and let $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ be the diameter. Assume that there exists a minimizer $x^{*}\in\operatorname{int}P$, i.e., there exists an $r>0$ with $B(x^{*},r)\subset P$. Consider the iterates of Algorithm 1 with $\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}},1\right\}$ or Algorithm 2. Then, if $\rho/\mu<1$, it holds that, for all $t\geq 1$,
where $c=\frac{\sqrt{2\mu}}{\mu-\rho}$, $M:=(L(1+\nu))^{1/\nu}$, and
$t_{0}:=\max\left\{\left\lfloor\log_{\frac{1}{1+\nu}\left(1+\frac{\nu\rho}{\mu}\right)}\left(\frac{(L(1+\nu)(1-\rho/\mu)^{\nu}c^{1+\nu}D^{2}/r^{1+\nu})^{2/(1-\nu)}}{LD^{2}}\right)\right\rfloor+2,1\right\}.$
Proof.
Let $h_{t}:=f(x_{t})-f^{*}$. Using Lemma 3.4, we have
@@
where the third inequality holds due to Proposition B.8 and the definition of $D$ and the last inequality holds due to the local PL inequality from Remark B.6 with $c=\frac{\sqrt{2\mu}}{\mu-\rho}$ and $M:=(L(1+\nu))^{1/\nu}$.
@@
$\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\geq L(1+\nu)D_{\phi}(v_{t},x_{t}),$ (D.6)
because of $11/\nu>$ 1. Using Lemma 2.4 and (D.6), we have
@@
where the third inequality holds because of (3.2) in Lemma 3.3, and the last inequality holds because of the local quadratic growth condition of $f$ with $\mu>0$. Therefore, we have
@@
When $\nu=1$, we have $h_{t}-h_{t+1}\geq\min\left\{\frac{1}{2}\left(1-\frac{\rho}{\mu}\right),\frac{r^{2}}{2Mc^{2}D^{2}}\right\}\cdot h_{t}$. This inequality and the initial bound $f(x_{1})-f^{*}\leq LD^{2}$ due to Lemma A.4 imply
@@
Using $f(x_{1})-f^{*}\leq LD^{2}$ and Lemma B.11 with $c_{0}=LD^{2}$, $c_{1}=\frac{\nu}{1+\nu}\left(1-\frac{\rho}{\mu}\right)$, $c_{2}=\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}$, and $\theta_{0}=\frac{1-\nu}{2\nu}>0$, we have the claim. ∎
@@
D.4 Proof of Theorem 5.3
In the same way as Theorem 4.4, the pyramidal width is the minimal $\delta>0$ satisfying Lemma B.9 (see also [Braun2022-ew, Lemma 2.26] or [Lacoste-Julien2015-gb, Theorem 3]). An upper bound exists on the primal gap for weakly convex functions.
Lemma D.6 (Upper bound on primal gap for weakly convex functions).
@@
$f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}\delta^{2}}\langle\nabla f(x),v^{\textup{A}}-v^{\textup{FW}}\rangle^{2}.$ (D.7)
@@
where the first inequality follows from $01-\rho/\mu>$ 0. Using (B.8) with $\psi=\nabla f(x)$ and $y=x^{*}$ and the above inequality, we obtain

### v4

@@ D.1 Sublinear Convergence
where the first inequality holds because of Lemma B.7 and the last inequality holds because of Lemma D.1. Subtracting $f^{*}$ on both sides, using $D_{\phi}(v_{t},x_{t})\leq D^{2}$, and rearranging leads to
@@
If $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu=1$ in Equation D.2, we have $f(x_{t})-f^{*}\leq 4LD^{2}/(t+2)$, which is the same as a sublinear convergence rate of the classical FW algorithm.
Remark D.3.
While the convergence rate Equation D.2 is the same as Vyguzov and Stonyakin (2025), Vyguzov and Stonyakin assume that the triangle scaling property holds for $D_{\phi}$. In contrast, we require significantly weaker assumptions: it is enough to assume that $\phi$ is strictly convex due to Lemma B.3 in order to establish Theorem D.2.
D.2 Proof of Theorem 4.2
Because Algorithm 2 is well-defined (see Remark 3.2), the convergence result of the FW algorithm using the adaptive step-size strategy (Algorithm 2) is essentially the same as the one that uses Bregman short steps (Theorems 4.2 and 5.3); up to small errors arising from the approximation whose precise analysis we skip for the sake of brevity.
Proof of Theorem 4.2.
Let $h_{t}:=f(x_{t})-f^{*}$. Using Lemma B.8, we have
@@
where the second inequality holds from $\langle f(x_{t}),x_{t}-v_{t}\rangle\geq r\|\nabla f(x_{t})\|$ in Proposition C.7, and the last inequality holds because $f$ satisfies the cc-Łojasiewicz inequality with $c=(q/\mu)^{1/q}$ (see Remark C.4), and $M:=(L(1+\nu))^{1/\nu}$.
(ii) $\gamma_{t}=1$: Using Equation D.1, we have
@@
This inequality and the initial bound $f(x_{1})-f^{*}\leq LD^{2}$ due to Lemma B.4 imply
@@
Using $f(x_{1})-f^{*}\leq LD^{2}$ and Lemma C.11 with $c_{0}=LD^{2}$, $c_{1}=\frac{\nu}{1+\nu}$, $c_{2}=\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}$, and $\theta_{0}=\frac{q-1-\nu}{\nu q}>0$ from $q>1+\nu$, we have the claim. ∎
D.3 Proof of Theorem 4.4
Proof of Theorem 4.4.
By using the induced guarantee on the primal gap via Lemma C.10, we have
$h_{t}=f(x_{t})-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{\langle\nabla f(x_{t}),v_{t}^{\textup{A}}-v_{t}^{\textup{FW}}\rangle}{\delta}\right)^{\frac{q}{q-1}}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{2\langle\nabla f(x_{t}),d_{t}\rangle}{\delta}\right)^{\frac{q}{q-1}},$ (D.3)
@@
where the first inequality holds due to Lemma B.8 with $\gamma_{t,\max}=1$ (Frank–Wolfe steps) and $\gamma_{t,\max}=\frac{\lambda_{v_{t}^{\textup{A}},t}}{1-\lambda_{v_{t}^{\textup{A}},t}}$ (away steps), and the second inequality holds because of Equation D.3 and $\langle\nabla f(x_{t}),d_{t}\rangle\geq\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\geq h_{t}$.
@@
In $q>1+\nu$, we have $h_{t}-h_{t+1}\geq\min\left\{\frac{\nu}{1+\nu},\frac{\nu}{1+\nu}\frac{((\mu/q)^{1/q}\delta/2)^{\frac{1+\nu}{\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}h_{t}^{\frac{q-1-\nu}{\nu q}}\right\}\cdot h_{t}$ for at least half of the iterations (Frank–Wolfe steps) and $h_{t+1}\leq h_{t}$ for the rest (away steps). The initial bound $f(x_{1})-f^{*}\leq LD^{2}$ holds generally for the Frank–Wolfe algorithm (see the proof of Theorem D.2 and (Braun et al., 2025, Remark 2.4)). We use Lemma C.11 with $c_{0}=LD^{2}$, $c_{1}=\frac{\nu}{1+\nu}$, $c_{2}=c_{1}\frac{((\mu/q)^{1/q}\delta/2)^{\frac{1+\nu}{\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}$, and $\theta_{0}=\frac{q-1-\nu}{\nu q}>0$ from $q>1+\nu$ and obtain the claim. ∎
Remark D.4 (Compatibility of parameters).
@@
$\frac{((\mu/q)^{1/q}\delta/2)^{\frac{1+\nu}{\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}=\left(\frac{\mu}{L}\right)^{1/\nu}\left(\frac{\delta}{D}\right)^{\frac{1+\nu}{\nu}}\frac{1}{2^{\frac{1+\nu}{\nu}}(1+\nu)^{2/\nu}D^{(1-\nu)/\nu}},$ (D.4)
which might be greater than 1 because $\frac{1}{D^{(1-\nu)/\nu}}\geq 1$ if $D<1$ and $\nu$ are small enough. In order to make Equation D.4 smaller than 1, $\nu$ should be 1, i.e., $(\nu,q)=(1,2)$. When $\nu=1$ and $q\neq 2$, it reduces to the $q>1+\nu$ case.
Appendix E Proof of Convergence Analysis for Nonconvex Optimization
E.1 Global Convergence
We show that Algorithm 1 with $\gamma_{t}=\gamma:=1/(T+1)^{\frac{1}{1+\nu}}$ globally converges to a stationary point, where $T\in\mathbb{N}$ is the number of iterations. Its proof is inspired by (Lacoste-Julien, 2016, Theorem 1) and (Pokutta, 2024, Theorem 4.7) and identical to those for the case when $\phi=\frac{1}{2}\|\cdot\|^{2}$.
Theorem E.1 (Global sublinear convergence for nonconvex optimization).
@@ E.1 Global Convergence
Substituting $x_{t+1}$ for $x_{+}$ and $x_{t}$ for $x$ in Lemma B.7, we have the following inequality:
@@
which is the same rate as (Lacoste-Julien, 2016, Theorem 1).
E.2 Local Sublinear Convergence
@@
Now we show sublinear convergence with $\gamma_{t}=\frac{2}{2+t}$. The proof is a modified version of Theorem D.2.
Theorem E.3 (Local sublinear convergence).
@@
Using Lemma B.7, we have
@@
where the last inequality holds because of Equation C.1 in Lemma C.2. Subtracting $f^{*}$ and rearranging provides
$f(x_{t+1})-f^{*}$ $\leq(1-\gamma_{t})(f(x_{t})-f^{*})+\frac{\rho\gamma_{t}}{2}\|x_{t}-x^{*}\|^{2}+L\gamma_{t}^{1+\nu}D_{\phi}(v_{t},x_{t})$ (E.3)
$\leq\left(1-\left(1-\frac{\rho}{\mu}\right)\gamma_{t}\right)(f(x_{t})-f^{*})+L\gamma_{t}^{1+\nu}D^{2},$ (E.4)
where the last inequality holds due to the quadratic growth condition. For $t=0$, using Equation E.3, we have
$f(x_{1})-f^{*}\leq\frac{\rho}{2}\|x_{0}-x^{*}\|^{2}+LD_{\phi}(v_{0},x_{0})\leq\left(\frac{\rho}{\sigma}+L\right)D^{2}\leq 2LD^{2}\leq\frac{2\mu}{\rho}LD^{2},$ (E.5)
where the second inequality holds because $\phi$ is $\sigma$-strongly convex and the last inequality holds because of $3<6/(2-\nu)\leq 2\mu/\rho$. Now we consider $t\geq 1$. Using Equation E.4 and $\gamma_{t}=\frac{2}{2+t}$, we have
@@
where the last inequality holds because of $(t+3)^{\nu}(t+3\rho/\mu)\leq(t+3)^{\nu}(t+2-\nu)\leq(t+2)^{1+\nu}$ from Equation E.1. ∎
@@
Note that Theorem E.3 does not require knowledge of the number of iterations $T$ ahead of time compared to Theorem E.1. We stress, nonetheless, that the latter can also be adjusted using a different step-size strategy to obtain an any-time guarantee; see Braun et al. (2025) for details for the standard Euclidean case, which can be generalized to our setup. Moreover, we can apply Theorem E.3 to the Euclidean FW algorithm, i.e., $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu=1$. Using Equation E.2, $D_{\textup{Euc}}:=\sup_{x,y\in P}\|x-y\|^{2}$, and $D_{\textup{Euc}}=\sqrt{2}D$, we obtain
@@
Theorem E.3 requires $\rho/\sigma\leq L$ and $\frac{3\rho}{\mu}\leq 2-\nu$. These assumptions are easy to satisfy.
Example E.4 (Example 3.1 in the arXiv version of Liao et al. (2024)).
@@
The function $f$ is not convex but $\rho$-weakly convex with $\rho=2$. A global optimal solution of $f$ is $x^{*}=-1$ and its value is $f(x^{*})=0$. Moreover, $f$ satisfies the quadratic growth condition with $0<\mu\leq 6$. Because $f$ is a quadratic function, we have $L\geq 6$. It holds that $2=\rho\leq L$ and $1=\frac{3\rho}{\mu}\leq 1<2-\nu$ (set $\mu=6$). Therefore, the assumption of Theorem E.3 holds.
In order to verify $\frac{3\rho}{\mu}\leq 2-\nu$, it is easier to examine the sufficient condition $\frac{3\rho}{\mu}\leq 1$ instead because the exact value of $\nu$ is difficult to estimate. On the other hand, without loss of generality, we can assume $\sigma=1$. When $\sigma>1$, $\rho/\sigma<\rho\leq L$ holds. When $\sigma<1$, we can use $\phi_{1}=\phi+\frac{1-\sigma}{2}\|\cdot\|^{2}$, which is 1-strongly convex. When $\phi$ is convex but not strongly convex, we can use $\phi_{2}=\phi+\frac{1}{2}\|\cdot\|^{2}$. Therefore, it suffices to verify $\rho/\sigma\leq\rho\leq L$, which often holds (for example, see an example of phase retrieval in Section 6.2).
E.3 Proof of Theorem 5.3
Theorems 5.35.4, and E.9 require the weak convexity and the local quadratic growth condition of $f$ with $\rho\leq\mu$. We show several examples as follows.
Example E.5.
We show some examples of $\rho$-weakly convex functions satisfying the (local) quadratic growth condition with $\rho<\mu$.
  1. (i)
    Let $(x)=\log(1+x^{2}). ff$ attains the minimum $f^{*}=0$ at $x^{*}=0$. The function $f$ is $1/4$-weakly convex because $f^{\prime\prime}(x)=-\frac{2(1-x^{2})}{(1+x^{2})^{2}}$ and its minimum is 1/4-1/4. Moreover, $(x-x^{*})^{2}=x^{2}\leq\frac{2}{\mu}\log(1+x^{2})=\frac{2}{\mu}(f(x)-f^{*})$ holds for any $x\in[f\leq\zeta]$. Letting $\mu=1/3$, we have $6\log(1+x^{2})-x^{2}\geq 0$ for any $x\in[f\leq 16]$. Note that $\zeta=16$ is not an exact value. In this case, we have $1/4=\rho<\mu=1/3$. Through simple calculations and visualization, we can obtain a larger value of $\mu$ than $1/3$. This function is used for image restoration (Boţ et al., 2016; Stella et al., 2017).
  2. (ii)
    Let $(x)=\frac{1}{2}x^{2}+2(1-e^{-x^{2}}). ff$ attains the minimum $f^{*}=0$ at $x^{*}=0$. The function $f$ is $(8e^{-3/2}-1)$-weakly convex because $f^{\prime\prime}(x)=1+4e^{-x^{2}}(1-2x^{2})$ and its minimum is $1-8e^{-3/2}\simeq-0.785$. Moreover, $(x-x^{*})=x^{2}\leq\frac{2}{\mu}(f(x)-f^{*})$ holds for any $x\in[f\leq\zeta]$. Letting $\mu=1$, we have $\frac{2}{\mu}(f(x)-f^{*})-x^{2}=4(1-e^{-x^{2}})\geq 0$ for any $x\in\mathbb{R}$. The function $f$ satisfies the global quadratic growth condition with $8e^{-3/2}-1=\rho<\mu=1$.
  3. (iii)
    Let $f$ be the function defined in Example E.4. We set $\mu=\mu_{0}$ for some $\mu_{0}\in(2,6]$ because $\mu$ can be from $(0,6]$. It holds that $2=\rho<\mu$.
Proof of Theorem 5.3.
Let $h_{t}:=f(x_{t})-f^{*}$. Using Lemma B.8, we have
@@
where the third inequality holds due to Proposition C.8 and the definition of $D$ and the last inequality holds due to the local PL inequality from Remark C.6 with $c=\frac{\sqrt{2\mu}}{\mu-\rho}$ and $M:=(L(1+\nu))^{1/\nu}$.
@@
$\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\geq L(1+\nu)D_{\phi}(v_{t},x_{t}),$ (E.6)
because of $1/\nu>1$. Using Lemma B.7 and Equation E.6, we have
@@
where the third inequality holds because of Equation C.1 in Lemma C.2, and the last inequality holds because of the local quadratic growth condition of $f$ with $\mu>0$. Therefore, we have
@@
When $\nu=1$, we have $h_{t}-h_{t+1}\geq\min\left\{\frac{1}{2}\left(1-\frac{\rho}{\mu}\right),\frac{r^{2}}{2Mc^{2}D^{2}}\right\}\cdot h_{t}$. This inequality and the initial bound $f(x_{1})-f^{*}\leq LD^{2}$ due to Lemma B.4 imply
@@
Using $f(x_{1})-f^{*}\leq LD^{2}$ and Lemma C.11 with $c_{0}=LD^{2}$, $c_{1}=\frac{\nu}{1+\nu}\left(1-\frac{\rho}{\mu}\right)$, $c_{2}=\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}$, and $\theta_{0}=\frac{1-\nu}{2\nu}>0$, we have the claim. ∎
@@
E.4 Proof of Theorem 5.4
In the same way as Theorem 4.4, the pyramidal width is the minimal $\delta>0$ satisfying Lemma C.9 (see also (Braun et al., 2025, Lemma 2.26) or (Lacoste-Julien and Jaggi, 2015, Theorem 3)). An upper bound exists on the primal gap for weakly convex functions.
Lemma E.6 (Upper bound on primal gap for weakly convex functions).
@@
$f(x)-f^{*}\leq\frac{2\mu}{(\mu-\rho)^{2}\delta^{2}}\langle\nabla f(x),v^{\textup{A}}-v^{\textup{FW}}\rangle^{2}.$
@@
where the first inequality follows from $1-\rho/\mu>0$. Using Equation C.8 with $\psi=\nabla f(x)$ and $y=x^{*}$ and the above inequality, we obtain